# Joint Estimation of Identity Verification and Relative Pose for Partial Fingerprints

### v1

@@ III-B Information Interaction
It should be noted that following two attention blocks have the same structure, including residual-connected Multi-Head Attention (MHA) and Feed-Forward Network (FFN) [70], only the data streams are differentiated by their respective expected effects. Specifically, let $\operatorname{SA}\left(f_{i},f_{j}\right)$ (Self-Attention) and $\operatorname{CA}\left(f_{i},f_{j}\right)$ (Cross-Attention) represent the corresponding sub-blocks in Fig. 3, the complete process of transformer block is defined as:
@@
The linear projection of $f_{i}$ and $f_{j}$ in MHA from left to right are $Query$, $Key$ and $Value$ matrices, which are then used for dot-product attention calculation formally as:
@@
where $d_{k}$ is the $Key$ dimention and the number of heads is fixed to $6$ at this stage. In order to further refine features while reducing information loss, a patch merging layer [71] is subsequently incorporated for downsampling. Finally, the positional encoding is applied again in the same way of Equation 1.
@@ III-C Identification Verification & Pose Prediction
$\left[\begin{array}[]{l}u_{\mathrm{a}}\\ v_{\mathrm{a}}\\ 1\end{array}\right]=\left[\begin{array}[]{ccc}\cos\theta&-\sin\theta&t_{\mathrm{x}}\\ \sin\theta&\cos\theta&t_{\mathrm{y}}\\ 0&0&1\end{array}\right]\cdot\left[\begin{array}[]{l}u_{\mathrm{b}}\\ v_{\mathrm{b}}\\ 1\end{array}\right]\;,$
@@ III-D1 Identity Verification Loss
For classification of identity verification, we adapt focal loss [72] to help our model focus on difficult samples:
@@
III-D2 Pose Prediction Loss
For regression of pose prediction, weighted mean square error is utilzed for supervision:
@@ IV Dataset Description
Comprehensive evaluations are conducted on extensive public datasets, including the representative benchmarks NIST SD14 [34], FVC2002 DB1_A & DB3_A [35], FVC2004 DB1_A & DB2_A [36], and FVC2006 DB2_A [37]. In addition, we also use a private partial fingerprint dataset THU Small to further examine the performance in smartphone situations. Table I presents a detailed description about the composition and usage of these datasets. In particular, we merge several fingerprint databases into a larger dataset, called the Hybrid Dataset (Hybrid DB) , to make the implementation of various experiments more convenient and sufficient. On the other hand, we further evaluate the performance of cross-domain generalization on other separate datasets. Rolled fingerprint is utilized because it covers a more complete area at different pressing poses than plain fingerprints. Image examples of different datasets are shown in Fig. 6. It can be seen that datasets selected in our experiments contains diverse fingerprint impressions with different modalities (normal, dry, wet, polluted, wrinkled, etc.).
TABLE I: All fingerprint datasets used in experiments
Type Dataset Scanners Technolohy Image Description Usage
Rolled NIST SD14 a[34] Inking Fingerprint Card $832\times 768$ 27,000 fingers $\times$ 2 impressions train & test c,h
Plain THU Small b Capacitive Smartphone $160\times 160$ 100 fingers $\times$ 8 impressions test d
FVC2002 DB1_A [35] Optical Identix $388\times 374$ 100 fingers $\times$ 8 impressions test e
FVC2002 DB3_A [35] Capacitive Precise Biometrics $300\times 300$ 100 fingers $\times$ 8 impressions test e
FVC2004 DB1_A [36] Optical CrossMatch $640\times 480$ 100 fingers $\times$ 8 impressions train & test f,h
FVC2004 DB2_A [36] Optical Digital Persona $328\times 364$ 100 fingers $\times$ 8 impressions train & test f,h
FVC2006 DB2_A [37] Optical BiometriKa $400\times 560$ 140 fingers $\times$ 12 impressions train & test g,h
@@
    Corresponding dataset is private.
@@
    A total of $108,000$ genuine match pairs and $108,000$ imposter match pairs are used in experiments.
  • d
    A total of $5,600$ genuine match pairs and $9,900$ imposter match pairs are used in experiments.
  • e
    A total of $5,600$ genuine match pairs and $5,600$ imposter match pairs are used in experiments.
  • f
    A total of $22,400$ genuine match pairs and $22,400$ imposter match pairs are used in experiments.
  • g
    A total of $73,920$ genuine match pairs and $73,920$ imposter match pairs are used in experiments.
  • h
@@ IV-A Parital Fingerprint Pair Simulation
Considering the lack of publicly available partial fingerprints, image patches of specified sizes are synthesized from corresponding rolled or plain datasets, following previous methods [8, 22]. For any two fingerprints from the same finger, we first use VeriFinger [10] to align them based on matching minutiae. Alignment in this way is accurate in most cases because selected original image pairs generally have large enough overlap areas of high quality. Subsequently, we calculated the common mask of aligned fingerprints and randomly select the patch center of one partial fingerprint from it. Taking this point as the center, another patch center is uniformly sampled in polar coordinates with a random angle ranging from $-180^{\circ}$ to $180^{\circ}$ and random radius ranging from 0 pixel to 100 pixels. Finally, full fingerprints are cropped into partial images of $160\times 160$, $128\times 128$ and $96\times 96$ with a random rotation from $[-180^{\circ},180^{\circ}]$ in our experiments, aiming to comprehensively evaluate the performance of representative methods under different patch sizes. The parameters of rigid translation and rotation are converted to ground truth $H$ according to the mapping relationship of Equation 6. It should be noted that each pair of fingerprints comes from two impressions respectively to avoid possible information leakage caused by the same grayscale relationship or distortion pattern.
@@ IV-B Matching Protocols
In order to balance the number of genuine / imposter matches and different fingerprint types / modalities, we implemented specific matching protocols during training and testing. For the FVC series, each fingerprint is combined with all mated fingerprints (with order) as genuine match pairs, and the first image of each pair is combined with a random fingerprint of another finger as a new imposter match pair. Taking FVC2002 DB1_A as an example, a total of $8\times 7\times 100\times 2=11,200$ image pairs are combined in this way. This protocol is also applied to NIST SD14 with a slightly different that symmetric matches are avoided, considering the large number of rolled images. All combined pairs of full fingerprints are then randomly cropped to simulate several partial fingerprint scenes, with the number of executions set to $1$ (FVC2002 DB1_A & DB3_A) and $4$ (the others). As mentioned above, some representative datasets are merged as Hybrid DB and subsequently divided, secifically $430,768$ for training and $22,672$ for testing, which are distinguished by suffixes ‘_A’ and ‘_B’ in the following experiments. The selection of genuine matches in real partial fingerprint dataset THU Small also follows this principle, while pairing the first impression of every finger with each other as imposter pairs. All detailed usage information is listed in corresponding footnotes of Table I.
@@ V Experiments
In this section, several representative state-of-the-art fingerprint algorithms are compared with our proposed approach sufficiently and comprehensively, including:
  • A-KAZE [7], a simple but comparative partial fingerprint recognition algorithm based on universal key points;
  • VeriFinger v12.0 [10], a widely used, top performing commercial software mainly based on minutiae;
  • DeepPrint [13], a global fixed-length representation based fingerprint recognition algorithm which is highly influential;
  • LDRF [51], another region representation based fingerprint recognition algorithm that extracts more flexible localized deep descriptors;
  • PFVNet [8], a state-of-the-art partial fingerprint verification algorithm.
In particular, since the effective information of partial fingerprints is significantly reduced compared to full-size fingerprints, the absolute pose rectification pre-step adopted in DeepPrint [13] and LDRF [51] cannot be performed as expected. Therefore, in this paper we use AlignNet, part of PFVNet [8], to estimate and correct the relative pose of image pairs as substitute, which is marked by subscript ‘*’.
We conducted extensive experiments to thoroughly evaluate the performance of above algorithms under partial fingerprints at different sizes, which in terms of score distribution, matching performance, alignment accuracy, efficiency and visual analysis. In addition, exhaustive ablation studies are demonstrated to validate the effectiveness of corresponding modules and strategies proposed in this paper.
@@ V-A Implementation Details
All partial fingerprint training processes are performed on the generated dataset Hybrid DB_A (where patch size is fixed at $160\times 160$) described in Section IV with a initial learning rate of $1\mathrm{e}{-3}$ (end of $1\mathrm{e}{-6}$), cosine annealing scheduler, default AdamW optimizer and batch size of $128$ until convergence (about $12$ epochs). The pre-training task proposed in Section III-A is trained for $300$ epochs using a subset of $5,000$ rolled images with corresponding random augmentation. On the other hand, considering that both fixed-length representation based methods [13, 51] require large amounts of training data, we completely reimplemented them on full-size fingerprints in the manner of [51]. Hybrid DB_B, which has the same source as training set, is used to test the performance under same data distribution. Besides, scenarios of other image sizes and datasets are directly used for evaluation without fine-tuning, which can also reflect the cross-domain adaptability of each model to a certain extent.
@@ V-B Score Distribution
We first present the score distribution of genuine and imposter matches on Hybrid DB_B with image size $160\times 160$ to qualitatively demonstrate the discriminative capabilities of different algorithms. Corresponding curves are shown in Fig. 7, where the scores of A-KAZE, VeriFinger, and LDRF are linearly mapped to a numerical range that approximates $[0,1]$ for intuitive comparison. A more discrete pairwise distribution (i.e., fewer overlapping areas) indicates that the corresponding algorithm can better distinguish whether a certain fingerprint pair comes from the same finger.
Moreover, three common indicators are further computed and listed in Table II to quantitatively measure the differences between genuine and imposter probability distributions. Among them, the Jensen–Shannon divergence (JS) is defined as:
@@
Figure 7: Probability density distribution of genuine and imposter matching scores on Hybrid DB_B with image size $160\times 160$. Scales on vertical axis are hidden because we are more concerned with the relative values.
TABLE II: Probability distribution distance of genuine and imposter scores on Hybrid DB_B with image size $160\times 160$.
@@
According to the results in Fig. 7 and Table II, our method outperforms all other methods, while VeriFinger [10] and PFVNet [8] each have sub-optimal performance in some parts. In particular, the distinguishing ability of LDRF* [51] is not ideal because it is designed based on the priori assumption of high-precision alignment (within about $9$ pixel and $4^{\circ}$ reported in original paper), which is difficult to satisfy in current partial fingerprint situations.
@@ V-C Matching Performance
TABLE III: Matching Performance with State-of-the-Art Algorithms and Our Approach on Four Representative Datasets of Partial Fingerprint in Three Sizes.
Size Sketch Method Hybrid DB_B a FVC2002 DB1_A a FVC2002 DB3_A a THU Small b
AUC AP TAR EER AUC AP TAR EER AUC AP TAR EER AUC AP TAR EER
160 $\times$ 160 Fig. LABEL:fig:intro_method_a A-KAZE [7] 0.70 0.77 0.28 0.37 0.86 0.90 0.56 0.22 0.77 0.83 0.36 0.31 0.87 0.88 0.68 0.21
VeriFinger [10] 0.88 0.91 0.56 0.18 0.97 0.98 0.85 0.07 0.93 0.95 0.76 0.12 0.91 0.91 0.76 0.15
Fig. LABEL:fig:intro_method_b DeepPrint* [13] 0.87 0.88 0.16 0.21 0.92 0.93 0.32 0.16 0.89 0.90 0.16 0.19 0.86 0.83 0.26 0.22
LDRF* [51] 0.64 0.65 0.02 0.40 0.78 0.80 0.10 0.29 0.72 0.75 0.07 0.34 0.71 0.66 0.13 0.35
Fig. LABEL:fig:intro_method_c PFVNet [8] 0.94 0.95 0.47 0.14 0.96 0.97 0.61 0.10 0.96 0.97 0.39 0.11 0.87 0.85 0.47 0.22
@@
128 $\times$ 128 Fig. LABEL:fig:intro_method_a A-KAZE [7] 0.63 0.68 0.15 0.42 0.73 0.80 0.32 0.34 0.67 0.72 0.14 0.39 0.74 0.74 0.40 0.34
VeriFinger [10] 0.74 0.79 0.29 0.33 0.85 0.89 0.55 0.22 0.79 0.83 0.43 0.29 0.80 0.80 0.52 0.28
Fig. LABEL:fig:intro_method_b DeepPrint* [13] 0.73 0.75 0.03 0.33 0.75 0.78 0.07 0.32 0.75 0.77 0.07 0.32 0.69 0.60 0.05 0.37
LDRF* [51] 0.59 0.60 0.01 0.44 0.63 0.65 0.02 0.41 0.63 0.65 0.03 0.41 0.60 0.50 0.03 0.43
Fig. LABEL:fig:intro_method_c PFVNet [8] 0.81 0.82 0.12 0.27 0.79 0.82 0.15 0.29 0.82 0.83 0.13 0.27 0.74 0.67 0.11 0.33
@@
96 $\times$ 96 Fig. LABEL:fig:intro_method_a A-KAZE [7] 0.52 0.53 0.02 0.49 0.54 0.58 0.05 0.49 0.52 0.54 0.01 0.49 0.53 0.43 0.06 0.48
VeriFinger [10] 0.60 0.63 0.09 0.41 0.66 0.70 0.19 0.38 0.63 0.66 0.12 0.39 0.67 0.62 0.24 0.37
Fig. LABEL:fig:intro_method_b DeepPrint* [13] 0.64 0.65 0.01 0.40 0.61 0.64 0.02 0.42 0.64 0.65 0.01 0.40 0.56 0.44 0.01 0.45
LDRF* [51] 0.52 0.53 0.00 0.49 0.55 0.56 0.01 0.47 0.54 0.55 0.00 0.47 0.53 0.42 0.01 0.48
Fig. LABEL:fig:intro_method_c PFVNet [8] 0.71 0.71 0.03 0.36 0.65 0.67 0.04 0.40 0.69 0.70 0.03 0.38 0.64 0.52 0.02 0.42
@@
  • a
    Image pairs are uniformly distributed within a relative translation of 0 to 100 pixels and relative rotation range of 360 degrees.
  • b
    Fingerprints are entered by volunteers in any comfortable pressing pose, with no special declaration or device restrictions.
Figure 8: Equal Error Rate matrix under different degrees of relative rigid transformation on Hybrid DB_B with image size $160\times 160$. Genuine matches are divided into 16 subsets based on the value of relative rotation $\theta$ and translation $T$, which in terms of Small, Medium, Large, and Huge segmented evenly by numerical range. Imposter matches are not modified again.
@@
  • Area Under Curve (AUC) based on Receiver Operating Characteristic (ROC) curves, to evaluate the overall performance in 1:1 verification;
  • Average-Precision (AP) based on Precision-Recall (PR) curves, to approximately evaluate the search performance under 1:N closed-set;
  • True Accept Rate (TAR) under specified False Accept Rate (FAR), to rigorously evaluate the identification performance 1:N open-set;
  • Equal Error Rate (EER), a comparable standard for system trade-offs.
@@
Table III presents the corresponding matching performance of state-of-the-art algorithms and our proposed approach on several representative datasets at different sizes. Furthermore, we group the genuine matches on Hybrid DB_B according to their relative rigid transformation relationship and compare these algorithms under different rotations and translations using EER. Fig. 8 shows the corresponding comparisons. Additionally, the Detection Error Tradeoff (DET) curves on all test databases are calculated in Fig. 9 to provide more complete information for analysis.
Figure 9: DET curves with state-of-the-art algorithms and our approach on partial fingerprints of $160\times 160$. Solid and dotted lines represent deep learning methods and traditional methods respectively.
@@
These encouraging experimental results strongly demonstrates the superiority of our method, which outperforms previous advanced methods in almost all cases. PFVNet [8] also shows remarkable performance, especially in indicators of AUC and AP. However, it can be observed that the relative performance deteriorates markedly when training and testing data cross domains. Region representation based methods[13, 51] are inferior to others, which is reasonable because the highly generalized form of fixed-length descriptors sacrifices many potential details for faster search speed. Specifically, although more flexible localized descriptors are designed in LDRF [51], their effectiveness is limited, as analyzed in Section V-B. On the other hand, the performance of all algorithms drops significantly as the image size decreases or distance increases, but our scheme still leads and exhibits attractive robustness and stability. It is worth noting that the impact of rotation is not evident in Fig. 8, because key points based matching score do not require pose rectification and the prediction error in other algorithms remains at a similar level across different relative angles (confirmed in the following experiments).
V-D Visual Analysis
Figure 10: Representative visualization results of genuine matches based on occlusion sensitivity. Red areas indicate a large influence on matching judgment, while blue areas indicate the opposite. White numbers on the upper left are the accuracy of corresponding methods when they are verified.
In order to provide more specific, vivid, and intuitive comparison with interpretability, we employ occlusion sensitivity [22] to approximate the contribution of distinct local areas to the overall matching judgment. Fig. 10 shows four representative visualization results. Query fingerprint in images pairs is centered and the search fingerprint is rigidly aligned based on the ground truth as the background. Subsequently, a $32\times 32$ mask is slid and applied to search fingerprints and the differences between scores before and after occlusion are calculated, which are normalized and overlaid as the heatmap of search fingerprint. Considering that there is no reference relationship between the matching scores directly obtained (or mapped) by different algorithms, we convert these scores into a comparable probability form and attached them in corresponding visualization results. Specifically, given the genuine matching score $s$ and overall distributions (shown in Fig. 7), the verification accuracy can be expressed as:
$P\left(y=1|s\right)$ (14)
where $P$ and $f$ represent the probability and probability density, $y$ represents the identity to be verified (1 is the same, 0 is different). It can be seen that algorithms based on key points and region representation expose their respective shortcomings when there is a lack of sufficient features in overlapping areas (row 1) and significant differences in ridge texture of non overlapping areas (row 2). PFVNet [8] has clear advantages in both cases, even with minimal overlap (row 3). However, it fails in the last scene with confusion in rotation. At the same time, our method demonstrates optimal stability and robustness in focusing on overlapping regions and gives correct decisions with the highest accuracy.
Meanwhile, some failure cases in Fig. 11 show that our method still needs to be improved in some extreme scenarios, such as: (i) sever defacement caused by incorrect collection conditions (row 1); (ii) significant misalignment of ridges caused by finger distortion (row 2); (iii) extremely weak texture, which refers to almost no minutiae or changes in orientation, as well as large modal differences (row 3).
V-E Alignment Accuracy

### v2

@@ III-B Interrelated Feature Extraction
It should be noted that the next two attention blocks have the same structure, including residual-connected Multi-Head Attention (MHA) and Feed-Forward Network (FFN) [50], only the data streams are differentiated by their respective expected effects. Specifically, let $\operatorname{SA}\left(f_{i},f_{j}\right)$ (Self-Attention) and $\operatorname{CA}\left(f_{i},f_{j}\right)$ (Cross-Attention) represent the corresponding sub-blocks in Fig. 3, the complete process of transformer block is defined as:
@@
The linear projection of $f_{i}$ and $f_{j}$ in MHA from left to right are $Query$, $Key$ and $Value$ matrices, which are then used for dot-product attention formally as:
@@
where $d_{k}$ is the $Key$ dimention and the number of heads is fixed to $6$ at this stage. In order to further condense features while reducing information loss, a patch merging layer [56] is subsequently incorporated for downsampling. Finally, the positional encoding is applied again in the same way of Equation 1.
@@ III-C Identification Verification & Pose Prediction
$\left[\begin{array}[]{l}u_{\mathrm{1}}\\ v_{\mathrm{1}}\\ 1\end{array}\right]=\left[\begin{array}[]{ccc}\cos\theta&-\sin\theta&t_{\mathrm{x}}\\ \sin\theta&\cos\theta&t_{\mathrm{y}}\\ 0&0&1\end{array}\right]\cdot\left[\begin{array}[]{l}u_{\mathrm{2}}\\ v_{\mathrm{2}}\\ 1\end{array}\right]\;,$
@@ III-D1 Identity Verification Loss
For classification of identity verification, we adapt focal loss [57] to help our model better focus on difficult samples:
@@
III-D2 Pose Alignment Loss
For regression of relate pose, weighted mean square error is utilzed for supervision:
@@ IV Dataset Description
Comprehensive evaluations are conducted on extensive public datasets, including the representative public benchmarks NIST SD14 [28], FVC2002 DB1_A & DB3_A [29], FVC2004 DB1_A & DB2_A [30], and FVC2006 DB2_A [31]. In addition, we also use a private partial fingerprint dataset THU Small to further examine the performance in capacitive smartphone situations. Table I presents a detailed description about the composition and usage of these datasets. In particular, we merge several fingerprint databases into a larger dataset, called the Hybrid Database (Hybrid DB) , to make the implementation of various experiments more convenient and sufficient. On the other hand, we further evaluate the performance of cross-domain generalization on other separate datasets. Rolled fingerprint is utilized because it covers a more complete area at different pressing poses than plain fingerprints. Image examples of different datasets are shown in Fig. 6. It can be seen that datasets selected in our experiments contains diverse fingerprint impressions with various modalities (normal, dry, wet, polluted, wrinkled, etc.).
TABLE I: All Fingerprint Datasets Used in Experiments.
Type Dataset Scanners Description Usage Genuine pairs Impostor pairs
Rolled NIST SD14 a[28] Inking 27,000 fingers $\times$ 2 impressions train & test c 108,000 108,000
Plain THU Small b Capacitive 100 fingers $\times$ 8 impressions test 5,600 9,900
FVC2002 DB1_A [29] Optical 100 fingers $\times$ 8 impressions test 5,600 5,600
FVC2002 DB3_A [29] Capacitive 100 fingers $\times$ 8 impressions test 5,600 5,600
FVC2004 DB1_A [30] Optical 100 fingers $\times$ 8 impressions train & test c 22,400 22,400
FVC2004 DB2_A [30] Optical 100 fingers $\times$ 8 impressions train & test c 22,400 22,400
FVC2006 DB2_A [31] Optical 140 fingers $\times$ 12 impressions train & test c 73,920 73,920
@@
    Corresponding dataset is private. Fingerprints are entered by volunteers in any comfortable pressing pose, with no special declaration or device restrictions. That is, there are quite a few genuine matching pairs with little or no overlap.
@@ IV-A Parital Fingerprint Pair Simulation
Considering the lack of publicly available datasets of partial fingerprints, image patches of specified sizes are synthesized from rolled or plain datasets, similar to previous works [7, 21]. For any two fingerprints from the same finger, we first use VeriFinger [9] to align them based on matching minutiae. Alignment in this way is accurate in most cases because original image pairs generally have large enough overlap areas of high quality. Subsequently, we calculated the common mask of aligned fingerprints and randomly select the patch center of one partial fingerprint from it. Taking this center as the zero point, another patch center is uniformly sampled in polar coordinates with a random angle ranging from $-180^{\circ}$ to $180^{\circ}$ and random radius ranging from 0 pixel to 100 / 70 / 20 pixels. Finally, full fingerprints are cropped into partial images of $160\times 160$, $128\times 128$ and $96\times 96$ with a random relative rotation from $[-180^{\circ},180^{\circ}]$, aiming to comprehensively evaluate the performance under different partial fingerprint sizes. The parameters of rigid translation and rotation are converted to ground truth $H$ according to Equation 6. It should be noted that each pair of partial fingerprints are cropped from different original images to avoid possible information leakage caused by the same grayscale relationship or distortion pattern.
@@ IV-B Matching Protocols
In order to balance the number of genuine / impostor matches and different fingerprint types / modalities, we implemented specific matching protocols during training and testing. For the FVC series, each fingerprint is combined with all mated fingerprints (with order) as genuine match pairs, and the first image of each pair is combined with a random impression of another finger as a new impostor match pair. Taking FVC2002 DB1_A as an example, a total of $8\times 7\times 100\times 2=11,200$ image pairs are combined in this way. This protocol is also applied to NIST SD14 with a slightly different that symmetric matches are avoided, considering the large number of rolled images. All combined pairs of full fingerprints are then randomly cropped to simulate several partial fingerprint scenes, with the number of executions set to $4$ (NIST SD14, FVC2004 DB1_A & DB2_A, FVC2006 DB2_A) for training  /  testing and $1$ (FVC2002 DB1_A & DB3_A) for testing. As mentioned above, some representative datasets are merged as Hybrid DB and subsequently divided, secifically $430,768$ for training and $22,672$ for testing, which are distinguished by suffixes ‘_A’ and ‘_B’ in the following experiments. The selection of genuine matches in real partial fingerprint dataset THU Small also follows this principle, while pairing the first impression of every finger with each other as impostor pairs. All detailed usage information is listed in corresponding footnotes of Table I.
@@ V Experiments
In this section, several representative state-of-the-art fingerprint algorithms are compared with our proposed approach, including:
  • A-KAZE [6], a simple but comparative partial fingerprint recognition algorithm based on commonly used key points;
  • VeriFinger SDK 12.0 [9], a widely used, top performing commercial software mainly based on minutiae;
  • DeepPrint [12], a global fixed-length representation extracted by deep network, which is highly influential;
  • DesNet [13], a descriptor network mainly designed for latent fingerprint, used to extract localized deep descriptors of densely sampled patches;
  • PFVNet [7], a state-of-the-art partial fingerprint verification algorithm recently proposed, which introduces multi-level features fusion and local self-attention mechanism.
In particular, since the effective information of partial fingerprints is significantly reduced compared to full-size fingerprints, the original pose rectification pre-step adopted in DeepPrint [12] and DesNet [13] cannot be performed as expected. Therefore, in this paper we use AlignNet, part of PFVNet [7], to estimate and correct the relative pose of image pairs as substitute, which is marked by subscript ‘*’.
We conducted extensive experiments to thoroughly evaluate the performance of above algorithms under partial fingerprints at different sizes, which in terms of score distribution, matching performance, alignment accuracy, visual analysis and efficiency. In addition, ablation studies are demonstrated to validate the effectiveness of corresponding modules and strategies proposed in this paper.
@@ V-A Implementation Details
All partial fingerprint training processes are performed on the generated dataset Hybrid DB_A with an initial learning rate of $1\mathrm{e}{-3}$ (end of $1\mathrm{e}{-6}$), cosine annealing scheduler, default AdamW optimizer and batch size of $128$ until convergence (about $12$ epochs). The pre-training task proposed in Section III-A is trained for $200$ epochs using a subset of $5,000$ rolled images with corresponding random augmentation. On the other hand, we completely reimplemented fixed-length representation based methods [12, 13] on full-size fingerprints in corresponding manner. Hybrid DB_B, which has the same source as training set, is used to evaluate the performance under same data distribution. Besides, scenarios of other image sizes and datasets are directly tested without fine-tuning to reflect the cross-domain adaptability.
@@ V-B Score Distribution
We first present the score distribution of genuine and impostor matches on Hybrid DB_B with image size $160\times 160$ to qualitatively demonstrate the discriminative capabilities of different algorithms. Corresponding curves are shown in Fig. 7, where the scores of A-KAZE [6], VeriFinger [9], and DesNet [13] are linearly mapped to appropriate ranges that approximate $[0,1]$ for intuitive comparison. A more separated pairwise distribution (i.e., fewer overlapping areas) indicates that the corresponding algorithm can better distinguish whether a certain fingerprint pair comes from the same finger.
Moreover, three common indicators are further computed and reported in Table II to quantitatively measure the differences between genuine and impostor probability distributions. Among them, the Jensen–Shannon (JS) divergence is defined as:
@@
Figure 7: Probability density distributions of genuine and impostor matching scores on Hybrid DB_B with image size $160\times 160$. Scales on vertical axis are hidden because we are more concerned with the relative values.
TABLE II: Probability Distribution Distance Between Genuine and Impostor Scores on Hybrid DB_B with Image Size $160\times 160$.
@@
According to the results in Fig. 7 and Table II, our method outperforms all other methods, while VeriFinger [9] and PFVNet [7] each have sub-optimal performance in some parts. In particular, the distinguishing ability of DesNet* [13] is not ideal because it is designed based on the prior assumption of high-precision alignment, which is difficult to satisfy in current partial fingerprint situations.
@@ V-C Matching Performance
TABLE III: Matching Performance of State-of-the-Art Algorithms and Our Approach on Four Representative Datasets of Partial Fingerprint with Three Different Sizes.
Size Sketch Method Hybrid DB_B FVC2002 DB1_A FVC2002 DB3_A THU Small
ACC AUC TAR EER ACC AUC TAR EER ACC AUC TAR EER ACC AUC TAR EER
160 $\times$ 160 Fig. LABEL:fig:intro_method_a A-KAZE [6] 0.67 0.70 0.28 0.37 0.82 0.86 0.56 0.22 0.73 0.77 0.36 0.31 0.89 0.87 0.68 0.21
VeriFinger [9] 0.85 0.88 0.56 0.18 0.95 0.97 0.85 0.07 0.90 0.93 0.76 0.12 0.92 0.91 0.76 0.15
Fig. LABEL:fig:intro_method_b DeepPrint* [12] 0.79 0.87 0.16 0.21 0.84 0.92 0.32 0.16 0.81 0.89 0.16 0.19 0.82 0.86 0.26 0.22
DesNet* [13] 0.74 0.81 0.05 0.27 0.79 0.87 0.16 0.21 0.76 0.84 0.08 0.25 0.77 0.78 0.20 0.29
Fig. LABEL:fig:intro_method_c PFVNet [7] 0.86 0.94 0.47 0.14 0.90 0.96 0.61 0.10 0.90 0.96 0.39 0.11 0.84 0.87 0.47 0.22
@@
128 $\times$ 128 Fig. LABEL:fig:intro_method_a A-KAZE [6] 0.60 0.63 0.22 0.39 0.69 0.73 0.45 0.27 0.63 0.67 0.20 0.34 0.79 0.74 0.40 0.34
VeriFinger [9] 0.72 0.74 0.39 0.27 0.82 0.85 0.71 0.13 0.77 0.79 0.55 0.20 0.83 0.80 0.51 0.28
Fig. LABEL:fig:intro_method_b DeepPrint* [12] 0.67 0.73 0.05 0.29 0.69 0.75 0.14 0.26 0.69 0.75 0.09 0.27 0.70 0.69 0.05 0.37
DesNet* [13] 0.64 0.69 0.02 0.37 0.66 0.71 0.04 0.35 0.64 0.69 0.02 0.36 0.70 0.67 0.06 0.38
Fig. LABEL:fig:intro_method_c PFVNet [7] 0.73 0.81 0.16 0.24 0.71 0.79 0.18 0.25 0.73 0.82 0.16 0.24 0.73 0.74 0.14 0.33
@@
96 $\times$ 96 Fig. LABEL:fig:intro_method_a A-KAZE [6] 0.52 0.52 0.07 0.49 0.54 0.54 0.26 0.45 0.52 0.52 0.07 0.48 0.66 0.53 0.06 0.47
VeriFinger [9] 0.59 0.60 0.25 0.31 0.65 0.66 0.53 0.20 0.62 0.63 0.26 0.24 0.74 0.67 0.25 0.37
Fig. LABEL:fig:intro_method_b DeepPrint* [12] 0.61 0.64 0.05 0.30 0.59 0.61 0.01 0.30 0.61 0.64 0.08 0.27 0.64 0.56 0.02 0.45
DesNet* [13] 0.58 0.60 0.01 0.42 0.57 0.60 0.01 0.43 0.56 0.58 0.01 0.44 0.65 0.58 0.01 0.45
Fig. LABEL:fig:intro_method_c PFVNet [7] 0.64 0.71 0.09 0.30 0.60 0.65 0.10 0.31 0.62 0.69 0.14 0.28 0.66 0.64 0.02 0.41
@@
Figure 8: Equal Error Rate matrix under different degrees of relative rigid transformation on Hybrid DB_B with image size $160\times 160$. Genuine matches are divided into 16 subsets based on the value of relative rotation $\theta$ and translation $T$, which in terms of Small, Medium, Large, and Huge segmented evenly by numerical range. Impostor matches are not modified again.
TABLE IV: Matching Performance of Different Score Level Fusion Strategies with image size $160\times 160$.
Dataset VeriFinger[9]     PFVNet[7]     PFVNet[7]     Proposed     Proposed Proposed
+ DeepPrint*[12] + VeriFinger[9] + DeepPrint*[12] + VeriFinger[9] + DeepPrint*[12] + PFVNet[7]
Hybrid DB_B      0.57 $\uparrow\!0.01$      0.66 $\uparrow\!0.19$      0.47 --      0.76 $\uparrow\!0.05$      0.74 $\uparrow\!0.03$      0.71 --
FVC2002 DB1_A      0.86 $\uparrow\!0.01$      0.87 $\uparrow\!0.26$      0.62 $\uparrow\!0.01$      0.94 $\uparrow\!0.01$      0.92 $\downarrow\!0.01$      0.95 $\uparrow\!0.02$
FVC2002 DB3_A      0.76 --      0.77 $\uparrow\!0.38$      0.49 $\uparrow\!0.10$      0.87 $\uparrow\!0.20$      0.68 $\uparrow\!0.01$      0.73 $\uparrow\!0.06$
THU Small      0.76 --      0.77 $\uparrow\!0.30$      0.47 --      0.80 $\uparrow\!0.10$      0.66 $\downarrow\!0.04$      0.66 $\downarrow\!0.04$
[…]
  • Scores of bold method are used as the baselines of corresponding columns. ‘+’ indicates another type of method for fusion.
@@
  • Accuracy (ACC) under optimal thresholds, to intuitively evaluate the correctness of model prediction;
@@
  • Area Under Curve (AUC) based on Receiver Operating Characteristic (ROC) curves, to approximately evaluate the overall classification performance;
[…]
  • Equal Error Rate (EER), where both false alarm and impostor pass errors are equal.
@@
Table III presents the corresponding matching performance of state-of-the-art algorithms and our proposed approach on several representative datasets of fingerprints with different sizes. Furthermore, we group the genuine matches on Hybrid DB_B according to their relative rigid transformation relationship and compare these algorithms under different rotations and translations using EER. Fig. 8 shows the corresponding comparisons. Additionally, the Detection Error Tradeoff (DET) curves on all test databases are calculated in Fig. .1 to provide more complete information for analysis.
@@
These encouraging experimental results strongly demonstrates the superiority of our proposed JIPNet, which outperforms previous advanced methods in almost all cases. PFVNet [7] also shows remarkable performance, especially in indicators of ACC and AUC. However, it can be observed that the relative performance deteriorates markedly when training and testing data cross domains. Region representation based methods[12, 13] are inferior to others, which is reasonable because the highly generalized form of fixed-length descriptors sacrifices many potential details for faster search speed. Specifically, although localized descriptors with deeper dimensions are designed in DesNet [13], their effectiveness is limited, as stated in Section V-B. On the other hand, the performance of all algorithms drops significantly as the image size decreases or distance increases, but our scheme still leads and exhibits attractive robustness and stability. It is worth noting that the impact of rotation is not evident in Fig. 8, because key points based matching score do not require pose rectification and the prediction error in other algorithms remains at a similar level across different relative angles (confirmed in the following experiments).
As analyzed above, these compared methods focus on different aspects of fingerprint attributes, exhibiting differentiated trends in score distribution and matching advantages. This phenomenon motivates us to conduct multiple fusion strategies to further observe their complementarity. Specifically, scores from two sets are linearly weighted and summed with the best classification accuracy. As represented in Table IV, most fusions effectively improves the corresponding performance, especially the introduction of minutiae information (VeriFinger) into other methods that rely on texture features, while our method still leads the way.
V-D Alignment Accuracy